# Massively multiplexed nanoscale magnetometry with diamond quantum sensors

### v1

@@ Acknowledgements.
Author contribution: KHC, JR, NPdL conceived the experiment platform. KHC and ZK built the optical setup with inputs from JR, BZ, and LN. KHC wrote the control software. BZ and JDT instructed the setup of the spatial light modulator. KHC and ZK collected and analyzed the data. KHC performed the numerical simulation and derivations. JDT and NPdL supervised the project. KHC, ZK, JR, and NPdL wrote the manuscript with input from all the authors.
@@
Figure 4: Multiplexed correlation measurements. a, Wide-field image showing the 4 targeted NV centers. The NV center spin and charge states are initialized by lasers A and B, where A addresses NV1-3 and B addresses only NV4, enabling selective reinitialization. b, Pulse sequences of the two types of correlation experiments carried out. We demonstrate the measurement of driven-spin correlations and spectroscopy of a correlated field using 4 NV centers. In both experiments, we prepare three NV centers (NV1-3) in the same spin state and one NV center (NV4) in the opposite spin state. In the driven-spin correlation experiment, we interleave MW pulses that rotate the spin by either $\theta$ or $\theta+\pi$ and sweep $\theta$. In the spectroscopy experiment, we run an XY8 sensing sequence with a fixed inter-pulse spacing while applying an AC magnetic field to the sample and sweeping the AC frequency. c, Driven-spin correlations between 4 NVs as a function of $\theta$. The three blue curves on the top show positive correlations between NV1-3, and the three red curves on the bottom show negative correlations between NV4 and NV1-3. d, Covariance noise spectroscopy of an applied field with 4 NV centers. The three NV pairs that start in the same spin state show positive correlations, while the three pairs that start in the opposite spin states show negative correlations.
@@ I Experimental details
An SLM (Santec SLM-210) is used to generate a laser pattern on the diamond sample to selectively address the NV centers chosen for SCC readout. We incorporate an open-source SLM control software package [19] for SLM calibration and phase pattern generation. The laser pattern is sampled from the main optical path onto a CMOS camera (Thorlabs CS165MU) for adaptive pattern calibration. The SLM pattern that corresponds to the locations of the selected NV centers is generated after selecting the NV positions, as shown in Fig. 1a.
@@ II Wide-field spin measurement of NV centers
where $\sigma_{0,1}^{2}$ and $\alpha_{0,1}$ are the variances and means, respectively, of the $m_{s}=0,1$ state photon number distributions [34]. The $\sigma_{\text{R}}$ of most NV centers are above 100, which is several times to an order of magnitude worse than the green readout using a confocal microscope (Extended Data Fig. 5). We attribute the high readout noise to the following three causes: (1) The EMCCD, unlike confocal microscopes, does not have a pinhole in the detection optical path to block out-of-plane light, so the EMCCD readout suffers from increased background fluorescence. (2) The spin contrast we obtain from the EMCCD using the protocol described above is lower than that obtained from a confocal microscope because when we integrate multiple pulsed experiments on one EMCCD exposure, we collect the fluorescence while the NV center spin is reinitialized, reducing the spin contrast. (3) The optical power density we use is approximately two orders of magnitude lower than the saturation power of a single NV center in a confocal microscope setup.
@@ VI Limits to multiplexed readout noise
The minimum $\sigma_{\text{R}}$ we have measured with our modified SCC scheme is approximately 2 to 3 times higher than observed in other works [36, 34]. By applying a rate equation model [40, 39], we study the dependence of $\sigma_{\text{R}}$ on ionization laser power and wavelength. In our simulation, we assume the following parameters of the NV center initial state: 70% NV- population, 95% initialization into $m_{s}$ = 0 state, and 1.6/6.7 photons on average for the NV0/NV- charge state readout, based on typical results of our charge state measurements. We evolve the system under the rate equation model with 95% $m_{s}$ = 0 or $m_{s}$ = 1 initial population that correspond to the cases where we apply no MW pulse or a $\pi-$pulse. At each time step, we extract the photon statistics of both cases from their instantaneous charge state distribution, $f_{0}(t)$ and $f_{\pi}(t)$:
@@
where $\textrm{Poisson}(\lambda)$ is the Poisson distribution with mean $\lambda$ and $p^{NV^{0,-}}_{0,\pi}$ denotes the NV center charge state population for either the neutral (0) or negative (-) charge state with the initial condition of applying a $\pi-$pulse ($\pi$) or not (0). As mentioned above, we assume $\lambda_{0}=1.6$ and $\lambda_{1}=6.7$ for the simulation. From the two photon readout distribution, we can then compute the corresponding statistics to extract the readout noise $\sigma_{\text{R}}$ with S3 to obtain $\sigma_{\text{R}}(t_{ion})$, where $t_{ion}$ is the evolution time that corresponds to the duration of the ionization pulse (Extended Data Fig. 6a). The model shows that for a given excitation laser power, as we increase the number of NV centers that we multiplex on (decrease the ionization power per NV center), the optimal ionization pulse time required to obtain the lowest readout noise at a given power increases (Extended Data Fig. 6b), which was also observed qualitatively in experiment. Using the same approach, we investigate the scaling of $\sigma_{\text{R}}$ as a function of number of multiplexed NV centers for different total available laser power (Extended Data Fig. 6c). With a lower total available laser power, $\sigma_{\text{R}}$ increases monotonically as the number of multiplexed NV centers increases. With a higher total power, $\sigma_{\text{R}}$ first decreases and then increases as the number of multiplexed NV centers increases. The scaling can be understood from the physical process behind SCC, where both too high or too low laser power decreases the ionization contrast between the $m_{s}=0$ and $m_{s}=\pm 1$ states. We also investigate the effect of the ionization laser wavelength on $\sigma_{\text{R}}$, which primarily determines the recombination rate from NV0 back to NV- [3], and obtain a lower floor for $\sigma_{\text{R}}$ when using a longer wavelength for ionization (Extended Data Fig. 6d).
@@ VIII Characterization of the NV centers for multiplexed covariance magnetometry
where $\tilde{\chi}_{i}(t)$ is the decoherence function, $\sigma_{\text{R}_{i}}$ is the spin readout noise, and $\phi_{C_{i}}$ is the phase accumulated by NV center $i$ due to a correlated field [34]. S11 shows that the measured magnetic field correlation is a function of both the NV readout noise and coherence. Since we are working with shallow, near-surface NV centers whose coherence and charge state stability could be corrupted by surface inhomogeneities, we characterize the NV properties before executing the multiplexed covariance experiment. To understand the coherence of the NV centers, we first run a wide-field XY8 dynamical decoupling sequence and sweep the interpulse spacing $\tau$ to characterize the NV coherence time $T_{2,XY8}$. The resulting $T_{2,XY8}$ map is shown in Extended Data Fig. 9. The wide-field $T_{2,XY8}$ data shows that the majority of the measured NV centers have coherence time over 15 µs, indicating that for the interpulse spacing $\tau$ = 250 ns used in our covariance magnetometry sensing sequence, the decoherence factor in correlation is $\sim 0.9^{2}$. Therefore, in the experiment, we pick NV centers that have coherence time greater than 15 µs. As for readout noise, the numbers that we measure from the selected NV centers for the multiplexed covariance magnetometry experiments are consistently around $\sigma_{\text{R}}\approx 12$-15 with minor fluctuations possibly due to the regeneration of laser spot pattern.

### v2

@@ Acknowledgements.
Author contribution: KHC, JR, NPdL conceived the experiment platform. KHC and ZK built the optical setup with inputs from JR, BZ, and LSN. KHC wrote the control software. BZ and JDT instructed the setup of the spatial light modulator. KHC and ZK collected and analyzed the data. KHC performed the numerical simulation and derivations. JDT and NPdL supervised the project. KHC, ZK, JR, and NPdL wrote the manuscript with input from all the authors.
@@
Correspondence and requests to materials should be addressed to N. P. de Leon.
@@
Figure 4: Multiplexed correlation measurements. a, Wide-field image showing the 4 targeted NV centers. The NV center spin and charge states are initialized by lasers A and B, where A addresses NV1-3 and B addresses only NV4, enabling selective reinitialization. b, Pulse sequences of the two types of correlation experiments carried out. We demonstrate the measurement of driven-spin correlations and spectroscopy of a correlated field using 4 NV centers. In both experiments, we prepare three NV centers (NV1-3) in the same spin state and one NV center (NV4) in the opposite spin state. In the driven-spin correlation experiment, we interleave MW pulses that rotate the spin by either $\theta$ or $\theta+\pi$ and sweep $\theta$. In the spectroscopy experiment, we run an XY8 sensing sequence with a fixed inter-pulse spacing while applying an AC magnetic field to the sample and sweeping the AC frequency. c, Driven-spin correlations between 4 NVs as a function of $\theta$. The three blue curves on the top show positive correlations between NV1-3, and the three red curves on the bottom show negative correlations between NV4 and NV1-3. The errorbars represent the standard error. d, Covariance noise spectroscopy of an applied field with 4 NV centers. The three NV pairs that start in the same spin state show positive correlations, while the three pairs that start in the opposite spin states show negative correlations.
@@ I Experimental details
An SLM (Santec SLM-210) is used to generate a laser pattern on the diamond sample to selectively address the NV centers chosen for SCC readout. We incorporate an open-source SLM control software package [kim_large-scale_2019] for SLM calibration and phase pattern generation. The laser pattern is sampled from the main optical path onto a CMOS camera (Thorlabs CS165MU) for adaptive pattern calibration. The SLM pattern that corresponds to the locations of the selected NV centers is generated after selecting the NV positions, as shown in Fig. 1a.
@@ II Wide-field spin measurement of NV centers
where $\sigma_{0,1}^{2}$ and $\alpha_{0,1}$ are the variances and means, respectively, of the $m_{s}=0,1$ state photon number distributions [rovny_nanoscale_2022]. The $\sigma_{\text{R}}$ of most NV centers are above 100, which is several times to an order of magnitude worse than the green readout using a confocal microscope (Extended Data Fig. 5). We attribute the high readout noise to the following three causes: (1) The EMCCD, unlike confocal microscopes, does not have a pinhole in the detection optical path to block out-of-plane light, so the EMCCD readout suffers from increased background fluorescence. (2) The spin contrast we obtain from the EMCCD using the protocol described above is lower than that obtained from a confocal microscope because when we integrate multiple pulsed experiments on one EMCCD exposure, we collect the fluorescence while the NV center spin is reinitialized, reducing the spin contrast. (3) The optical power density we use is approximately two orders of magnitude lower than the saturation power of a single NV center in a confocal microscope setup.
@@ VI Limits to multiplexed readout noise
The minimum $\sigma_{\text{R}}$ we have measured with our modified SCC scheme is approximately 2 to 3 times higher than observed in other works [shields_efficient_2015, rovny_nanoscale_2022]. By applying a rate equation model [yuan_charge_2020, wirtitsch_exploiting_2023], we study the dependence of $\sigma_{\text{R}}$ on ionization laser power and wavelength. In our simulation, we assume the following parameters of the NV center initial state: 70% NV- population, 95% initialization into $m_{s}$ = 0 state, and 1.6/6.7 photons on average for the NV0/NV- charge state readout, based on typical results of our charge state measurements. We evolve the system under the rate equation model with 95% $m_{s}$ = 0 or $m_{s}$ = 1 initial population that correspond to the cases where we apply no MW pulse or a $\pi-$pulse. At each time step, we extract the photon statistics of both cases from their instantaneous charge state distribution, $f_{0}(t)$ and $f_{\pi}(t)$:
@@
where $\textrm{Poisson}(\lambda)$ is the Poisson distribution with mean $\lambda$ and $p^{NV^{0,-}}_{0,\pi}$ denotes the NV center charge state population for either the neutral (0) or negative (-) charge state with the initial condition of applying a $\pi-$pulse ($\pi$) or not (0). As mentioned above, we assume $\lambda_{0}=1.6$ and $\lambda_{1}=6.7$ for the simulation. From the two photon readout distribution, we can then compute the corresponding statistics to extract the readout noise $\sigma_{\text{R}}$ with S3 to obtain $\sigma_{\text{R}}(t_{ion})$, where $t_{ion}$ is the evolution time that corresponds to the duration of the ionization pulse (Extended Data Fig. 6a). The model shows that for a given excitation laser power, as we increase the number of NV centers that we multiplex on (decrease the ionization power per NV center), the optimal ionization pulse time required to obtain the lowest readout noise at a given power increases (Extended Data Fig. 6b), which was also observed qualitatively in experiment. Using the same approach, we investigate the scaling of $\sigma_{\text{R}}$ as a function of number of multiplexed NV centers for different total available laser power (Extended Data Fig. 6c). With a lower total available laser power, $\sigma_{\text{R}}$ increases monotonically as the number of multiplexed NV centers increases. With a higher total power, $\sigma_{\text{R}}$ first decreases and then increases as the number of multiplexed NV centers increases. The scaling can be understood from the physical process behind SCC, where both too high or too low laser power decreases the ionization contrast between the $m_{s}=0$ and $m_{s}=\pm 1$ states. We also investigate the effect of the ionization laser wavelength on $\sigma_{\text{R}}$, which primarily determines the recombination rate from NV0 back to NV- [aslam_photo-induced_2013], and obtain a lower floor for $\sigma_{\text{R}}$ when using a longer wavelength for ionization (Extended Data Fig. 6d).
@@ VIII Characterization of the NV centers for multiplexed covariance magnetometry
where $\tilde{\chi}_{i}(t)$ is the decoherence function, $\sigma_{\text{R}_{i}}$ is the spin readout noise, and $\phi_{C_{i}}$ is the phase accumulated by NV center $i$ due to a correlated field [rovny_nanoscale_2022]. S11 shows that the measured magnetic field correlation is a function of both the NV readout noise and coherence. Since we are working with shallow, near-surface NV centers whose coherence and charge state stability could be corrupted by surface inhomogeneities, we characterize the NV properties before executing the multiplexed covariance experiment. To understand the coherence of the NV centers, we first run a wide-field XY8 dynamical decoupling sequence and sweep the interpulse spacing $\tau$ to characterize the NV coherence time $T_{2,XY8}$. The resulting $T_{2,XY8}$ map is shown in Extended Data Fig. 9. The wide-field $T_{2,XY8}$ data shows that the majority of the measured NV centers have coherence time over 15 µs, indicating that for the interpulse spacing $\tau$ = 250 ns used in our covariance magnetometry sensing sequence, the decoherence factor in correlation is $\sim 0.9^{2}$. Therefore, in the experiment, we pick NV centers that have coherence time greater than 15 µs. As for readout noise, the numbers that we measure from the selected NV centers for the multiplexed covariance magnetometry experiments are consistently around $\sigma_{\text{R}}\approx 12$-15 with minor fluctuations possibly due to the regeneration of laser spot pattern.